# Wandering around: A bioinspired approach to visual attention through object motion sensitivity

### v1

@@ 2 Related work
The turning point in advancing selective attention for object detection was marked by the integration of ‘Gestalt laws’ into bottom-up saliency mechanisms [russell2014model], ensuring the perceptual grouping of closed contours as potential objects within the scene. The frame-based proto-object implementation [russell2014model] has been further adapted for event-based cameras [iacono2019proto] for the humanoid robot iCub reducing computational load by eliminating the initial processing stage of edge extraction achieved by utilising the inherent capabilities of the event-based cameras, which interpret the scene with leading and trailing edges, along with polarity, and therefore can be assumed to represent an edge map of the scene. The system ignores clutter and non-proto-object shapes while providing an online saliency map approximately every $\sim$100 ms. A sparse event-based depth implementation of the model [ghosh2022event] has shown a more localised response over frame-based implementation [hu2016proto] ensuring a task-dependent response prioritising proto-objects closer to the observer generating the event-based disparity map and the saliency map approximately every $\sim$170 ms. The final spiking-based implementation of the model [Dangelo_etal22], directly deployed on the neuromorphic SpiNNaker platform, achieves a significant reduction in latency (approximately $\sim$16 ms), although it requires an unmanageable number of platforms (6 SpiNNaker boards) to run the full implementation across different orientations and scales of the von Mises filters.
@@ 3 Methods
Figure 1 provides an overview of the entire architecture. The Speck device is mounted on a FLIR pan-tilt robotic unit (VERSION: 3.02). A stream of events is generated by the DVS integrated into the Speck device. These events are directly processed by the spiking Object Motion Sensitivity (sOMS) model, which produces the OMS map as the final output, enhancing object motion. The OMS map is subsequently fed into the SNN Proto-object model to detect proto-objects and identify the most salient point in the scene. The pixel coordinates of the maximum salient point $P(x,y)$ are then input into the SNN Attention Control (sAC) population, which generates the appropriate pan and tilt commands to foveate on the salient point. The system then performs fixational eye movements generating events to detect the next salient point. The simulation is executed on macOS 15.1.1 with an M2 chip, utilizing the Metal Performance Shaders (MPS) device. The system employs the sinabs library from SynSense for visual processing and Nengo [bekolay2014nengo] [^3] https://www.nengo.ai for the control mechanism. To facilitate future deployment on neuromorphic hardware and bridge the gap between simulation and real-world implementation, the Speck device is mounted on the PTU, using only events from the DVS sensors. This setup serves as an intermediate step toward full deployment, enabling us to demonstrate the architecture’s capabilities.
@@ 3.1 Spiking Object Motion Sensitivity (sOMS) model
The sOMS model receives input from the events generated by the DVS on the Speck device, which has a 128$\times$128 pixel resolution. Central to this implementation are the OMS kernels inspired by the approach outlined in [clerico2024retina], which are implemented as two Gaussian distributions (Equation 1), using kernels of 8 pixels (see Figure 2) and different $\sigma$ for the centre (c) and surround (s) kernels. Unlike the implementation proposed by Clerico et al. [clerico2024retina], the model simulates the behaviour of OMS cells using the sinabs library from SynSense.
Figure 2: View of the (a) center and (b) surround Gaussian kernels of the Object Motion Sensitivity (OMS) model.
$G(x,\mu,\sigma)=\frac{1}{(2\pi)^{d/2}|\sigma|^{1/2}}\exp\left(-\frac{1}{2}(x-\mu)^{\top}\sigma^{-1}(x-\mu)\right)$ (1)
Where $\mu_{c}$ = (0,0), which is equal to $\mu_{s}$; $\sigma_{c}$ = 1 and $\sigma_{s}$ = 4 (see Figure 2). The model represents the center and surround Gaussian distributions through their respective weights. This is achieved via a Leaky Integrate-and-Fire (LIF) Spiking Convolutional Neural Network (sCNN) from the SynSense library, sinabs, with $\tau$ set to 0.02 seconds. The events collected are then fed into the sOMS model, which collects events with an empirical update rate of 20 ms. This update can be reduced to a minimum of 1 ms when running events in real-time for the online demonstrator. The collection of events assumed to represent the edge map of the scene, Event map, is processed through the center and surround networks. The resulting center map is then subtracted from the surround map (Equation 2). The final OMS map is then obtained from the initial Event map as described by Clerico et al. [clerico2024retina] (Equation 2).
$OMS_{map}=Events_{map}[(center(Events_{map})-surround(Events_{map}))>\alpha]$ (2)
@@ 3.1.1 Characterisation of the sOMS model
The characterisation of the sOMS has been inspired by the recordings of real cells conducted by Ölveczky et al. [olveczky2003segregation], where the activity of cells was measured under various conditions: Eye+object, Eye only and Object only. The stimuli consist of moving gratings, with both the background and central region in motion. In the case of Eye+object, both the background and the central region move at different speeds; in Eye only, only the background is moving; and in Object only, only the central part is moving (see Figure 3).
Experiment Parameters RGB stimuli Event map OMS map
1 Eye+object: Background (sf=0.3; s=0.01) Foreground (sf=3; s=0.09) Refer to caption Refer to caption Refer to caption
2 Eye only: Background (sf=0.3; s=0.01) Refer to caption Refer to caption Refer to caption
3 Object only: Foreground (sf=3; s=0.09) Refer to caption Refer to caption Refer to caption
@@
Figure 3: From left to right: Experiment, Parameters, RGB stimuli, Event map and OMS map in three different situations: Eye+object, Eye only and Object only. Where ${fs}$ is the spatial frequency and ${s}$ is the speed in cycles per frame of the moving grating. The figure shows the sOMS model enhancing object motion, where white and black represent positive and negative polarities on the Event map, and white indicates the spikes of the sOMS model on the sOMS map.
The stimuli have been generated using the PsychoPy library [peirce2007psychopy]. The gratings for both the background and the central object move with a speed $s$ and a spatial frequency $fs$. The stimuli presented to the model are RGB video clips with a duration of $4s$ at $60$ fps. The stimuli are then converted into events using the ICNS Event-Based Camera Simulator (IEBCS) [^4] https://github.com/neuromorphicsystems/IEBCS library using the following standard parameters of the simulator: ${dt}=10000$, ${lat}=100$, ${jit}=10$, ${ref}=100$, ${tau}=300$, ${th}=0.1$ and $th_{noise}=0.01$. Where ${dt}$ is the delta time, ${lat}$ is the latency, ${jit}$ the jitter, ${ref}$ the refractory period, ${tau}$ the time constant, ${th}$ the positive/negative events generation threshold, and $th_{noise}$ is the added noise sampled from two distributions acquired using a real sensor under 161 lux. Figure 3 illustrates an example of the RGB stimuli, the converted event frame, and the OMS maps in the three different cases: Eye+Object, Eye Only and Object Only. The sOMS model encodes object motion sensitivity [clerico2024retina, olveczky2003segregation]: it produces a robust response when the central object moves against background motion in the Eye+Object scenario and, as expected, in the Object Only scenario.
Refer to caption Refer to caption
(a) (b)
@@
Figure 4: Average Mean firing rate (MFR) and mean inter-spike interval (ISI) for the Eye+Object case under two scenarios: (a) varying center ($\sigma_{c}$) and surround ($\sigma_{s}$) values of the kernels (pixels) [$\sigma_{c}=1,\sigma_{s}=4$; $\sigma_{c}=2,\sigma_{s}=4$; $\sigma_{c}=3,\sigma_{s}=4$; $\sigma_{c}=4,\sigma_{s}=4$; $\sigma_{c}=2,\sigma_{s}=8$; $\sigma_{c}=4,\sigma_{s}=8$] (Figure 2), and (b) varying both $\sigma$ values and kernel sizes [$\sigma_{c}=1,\sigma_{s}=4$, s=8; $\sigma_{c}=4,\sigma_{s}=8$, s=16; $\sigma_{c}=8,\sigma_{s}=16$, s=32].
@@
The Gaussian kernel parameters and sizes were analysed based on their responses in the Eye+Object scenario. Figure 4a) illustrates how decreasing the ratio between $\sigma_{c}$ and $\sigma_{s}$ proportionally reduces both the mean firing rate (MFR) and the mean inter-spike interval (ISI) of the system, attributed to the diminished inhibitory effect of the surround. The final two edge cases, where $\sigma_{s}$ matches the kernel size, exhibit a slight reduction in both metrics, leading to less pronounced object motion contrast enhancement. Larger kernel sizes, as shown in Figure 4 b), confirm the same response trend, with increased activity observed for the largest size. This effect is attributed to the smoothing influence of the Gaussian kernels on the image, resulting in a larger portion of the visual field surpassing the threshold ($\alpha$). An increased MFR would lead to heightened system activity, resulting in unnecessary power consumption, that in neurons chips scales with the number of spikes, as the response is already functional for $\sigma_{c}=1$, $\sigma_{s}=4$, and a kernel size, $s_{k}$, of 8 pixels. This analysis, therefore, validates the parameters proposed in [clerico2024retina].
@@
To demonstrate the scale invariance of the model, the characterisation also includes different spatial frequencies of the stimuli in the Eye+Object scenario. Figure 5 shows an example of the RGB stimuli, the converted event frame, and the OMS maps with different values of spatial frequency, ${fs}$, and the speeds of a cycle on each iteration of the loop, ${s}$, for the object (Foreground) and the Background. As shown by the OMS maps in the qualitative assessment in Figure 5, the response of the sOMS model effectively enhances object motion in all cases.
Experiment Parameters RGB stimuli Event map OMS map
4 Background (sf=0.2; s=0.01) Foreground (sf=3; s=0.09) Refer to caption Refer to caption Refer to caption
5 Background (sf=1; s=0.01) Foreground (sf=3; s=0.09) Refer to caption Refer to caption Refer to caption
6 Background (sf=4; s=0.01) Foreground (sf=3; s=0.09) Refer to caption Refer to caption Refer to caption
7 Background (sf=3; s=0.01) Foreground (sf=0.2; s=0.09) Refer to caption Refer to caption Refer to caption
8 Background (sf=3; s=0.01) Foreground (sf=1; s=0.09) Refer to caption Refer to caption Refer to caption
9 Background (sf=3; s=0.01) Foreground (sf=4; s=0.09) Refer to caption Refer to caption Refer to caption
@@
Figure 5: From left to right: Experiment, Parameters, RGB stimuli, Event map and OMS map for the Eye+object case with different spatial frequencies of the background and the foreground. Where ${fs}$ is the spatial frequency and ${s}$ is the speed in cycles per frame of the moving grating. The colors shown on the maps are the same as those in Figure 3.
@@
Further investigations analyze the average MFR and mean ISI of the population activity (see Figure 6) across all experiments presented in Figure 3 and Figure 5, as these metrics provide complementary insights into the dynamics of neural activity. The average MFR captures the overall firing rate, offering a measure of activity intensity, while the mean ISI reflects the temporal structure of spiking, highlighting patterns in neural coordination and timing. The responses confirm the periodicity of the moving gratings, with similar MFR and mean ISI observed across all scenarios, except for Experiment 2 (Eye only). In this case, the OMS model remains effective in suppressing most background activity, even without object motion, which is typically required for its optimal functioning. This results in a higher mean MFR compared to the other cases and a higher mean ISI, confirming the absence of an object in motion in the scene. The standard deviation of the mean MFR highlights the responsiveness of only a portion of the visual field in all cases. In contrast, the standard deviation of the mean ISI approaches infinity due to the majority of neurons spiking only once during the simulation.
@@
Figure 6: Average MFR and mean ISI for all experiments shown in Figure 3 and Figure 5. Experiments 1, 2, and 3 correspond to Eye+Object, Eye only, and Object only, respectively, while experiments 4 through 9 represent Eye+Object scenarios with varying spatial frequencies (${fs}$) and speeds (${s}$) in cycles per iteration, as depicted in Figure 5.
@@
Parameters 10 11 12 13 14 15 16
Eye+object: Background (sf=0.3) Foreground (sf=3)
$\tau$ = 0.1 s
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$s_{B}$ = 0.01 $s_{F}$ = 0.01 $s_{B}$ = 0.03 $s_{F}$ = 0.01 $s_{B}$ = 0.05 $s_{F}$ = 0.01 $s_{B}$ = 0.09 $s_{F}$ = 0.01 $s_{B}$ = 0.05 $s_{F}$ = 0.05 $s_{B}$ = 0.09 $s_{F}$ = 0.05 $s_{B}$ = 0.13 $s_{F}$ = 0.05
Figure 7: Example of OMS map for different speeds (where $s_{F}\leq s_{B}$) for the background ($s_{B}$) and foreground ($s_{F}$) in the case of the Eye+object scenario for the Background ($fs$=0.3) and the Foreground ($fs$=3).
In all cases examined (Experiments 10–16, Figure 7), the sOMS model appears unable to enhance object motion effectively (same applies for $\tau$ = 0.01), resulting in poor performance in object motion segmentation when egomotion-induced events are generated from a background moving faster than the foreground. The failure of motion detection can be attributed to two critical mechanisms. First, spatiotemporal integration between center and surround regions becomes ineffective when object-induced events are temporally sparse or significantly lower than background-induced events. Under such conditions, the membrane potential fails to reach the threshold required for motion detection or the dominance of background activity suppresses the differential signal needed to distinguish motion.
Second, from an active vision perspective, the motion dynamics of the visual sensor critically impact detection reliability. The challenge of motion detection under rapid camera movement is a well-documented issue in optic flow estimation [pauwels2006optic]. When camera movement exceeds a specific velocity threshold, local motion contrast becomes unresolvable, directly compromising the fundamental assumptions of optical flow estimation.
@@ 3.2 SNN Proto-Object model
This implementation builds upon the event-driven, saliency-based proto-object model previously developed for GPU [iacono2019proto] and SpiNNaker [Dangelo_etal22]. The primary difference from the earlier implementations is the use of the LIF sCNN sinabs layer, with weights represented by the von Mises kernel distribution (Equation 3).
$VM_{\theta}(x,y)=\frac{\exp(\rho\cdot R_{0}\cdot\cos(tan^{-1}(-y,x)-\theta)}{I_{0}(\sqrt{x^{2}+y^{2}-R_{0}})}$ (3)
@@
$G1=\sum_{\theta}B1_{\theta}*VM_{\theta}$ $G2=\sum_{\theta}B2_{\theta}*VM_{\theta+\pi}$ $G1^{*}=\sum_{\theta}B1_{\theta}*VM_{\theta+\pi}$ $G2^{*}=\sum_{\theta}B2_{\theta}*VM_{\theta}$ (5)
These orientations and scales ($scales=3$) are pooled (Equation 5) to produce the final grouping response, which generates the saliency map of the scene and identifies the $(x,y)$ coordinates of the most salient point (Equation  6 and Figure 1).
$P(x,y)=\arg\max_{(x,y)}\big{(}(G1-G1^{*})+(G2-G2^{*})\big{)}.$ (6)

### v2

@@ 2 Related work
The turning point in advancing selective attention for object detection was marked by the integration of ‘Gestalt laws’ into bottom-up saliency mechanisms [15], ensuring the perceptual grouping of closed contours as potential objects within the scene. The frame-based proto-object implementation [15] has been further adapted for event-based cameras [30] for the humanoid robot iCub reducing computational load by eliminating the initial processing stage of edge extraction achieved by utilising the inherent capabilities of the event-based cameras, which interpret the scene with leading and trailing edges, along with polarity, and therefore can be assumed to represent an edge map of the scene. The system ignores clutter and non-proto-object shapes while providing an online saliency map approximately every $\sim$100 ms. A sparse event-based depth implementation of the model [31] has shown a more localised response over frame-based implementation [20] ensuring a task-dependent response prioritising proto-objects closer to the observer generating the event-based disparity map and the saliency map approximately every $\sim$170 ms. The final spiking-based implementation of the model [39], directly deployed on the neuromorphic SpiNNaker platform, achieves a significant reduction in latency (approximately $\sim$16 ms), although it requires an unmanageable number of platforms (6 SpiNNaker boards) to run the full implementation across different orientations and scales of the von Mises filters.
@@ 3 Methods
Figure 1 provides an overview of the entire architecture. The Speck device is mounted on a FLIR pan-tilt robotic unit (VERSION: 3.02). A stream of events is generated by the DVS from the Speck device. These events are directly processed by the spiking Object Motion Sensitivity (sOMS) model, which produces the OMS map as the final output, enhancing object motion. The OMS map is subsequently fed into the SNN Proto-object model to detect proto-objects and identify the most salient point in the scene. The pixel coordinates of the maximum salient point $P(x,y)$ are then input into the SNN Attention Control (sAC) population, which generates the appropriate pan and tilt commands to foveate on the salient point. The system then performs fixational eye movements generating events to detect the next salient point. The simulation is executed on macOS 15.1.1 with an M2 chip, utilizing the Metal Performance Shaders (MPS) device. The system employs the sinabs library from SynSense for visual processing and Nengo [81] [^3] https://www.nengo.ai for the control mechanism. To facilitate future deployment on neuromorphic hardware and bridge the gap between simulation and real-world implementation, the Speck device is mounted on the PTU, using only events from the DVS sensors. This setup serves as an intermediate step toward full deployment, enabling us to demonstrate the architecture’s capabilities.
@@ 3.1 Spiking Object Motion Sensitivity (sOMS) model
The sOMS model receives input from the events generated by the DVS on the Speck device, which has a $\times128$ pixel resolution. Central to this implementation are the OMS kernels inspired by the approach outlined in [52], which are implemented as two Gaussian distributions, using kernels of 8 pixels (see Figure 2) and different $\sigma$ for the centre (c) and surround (s) kernels. Unlike the implementation proposed by Clerico et al. [52], the model simulates the behaviour of OMS cells using the sinabs library from SynSense.
Figure 2: From left to right, an abstract representation of the networks to generate the OMS map. The Event map is processed through the center and surround networks consisting of a convolutional (Conv) layer followed by a LIF layer, producing the intermediate maps used to generate the OMS map, as shown in Equation 1.
The model employs center (c) and surround (s) Gaussian distributions, both centered at $\mu_{c}=\mu_{s}=(0,0)$, with the standard deviation $\sigma_{c}=1$ and $\sigma_{s}=4$ (see Figure 2). These distributions are implemented through their respective weights through a Leaky Integrate-and-Fire (LIF) Spiking Convolutional Neural Network (sCNN) from the SynSense library, sinabs. The network consists of two parallel pathways, center and surround. Each pathway includes a convolutional layer with a Gaussian kernel, followed by a layer of LIF neurons with a membrane time constant $\tau=0.02s$ (see Figure 2). Motion sensitivity in the model arises from the temporal integration properties of LIF neurons, which enable responses to spatio-temporal activation patterns, facilitating the detection of local contrastive motion. The events are collected and then fed into the sOMS model, which updates at an empirical rate of $20ms$. This update rate can be reduced to a minimum of $1ms$ for real-time processing in the online demonstrator. The collected events, assumed to represent the edge map of the scene (the Event map), are processed through the center and surround kernels. The resulting center map is then subtracted from the surround map (Equation 1). The final OMS maps for the positive and negative polarities are then obtained similarly to Clerico et al. [52], as the thresholded version of the initial Event map (Equation 1). For simplicity, throughout the remainder of this paper, we refer to their sum, i.e., the combined positive and negative polarity maps, as the OMS map. However, unlike their approach, we retain the signed difference between the center and surround filters to preserve the center-surround structure, which would be lost if we used the absolute value.
$OMS_{map}$ $=\text{center}(Events_{map})-\text{surround}(Events_{map})$ (1)
$OMS_{map}$ $=Events_{map}\left[1-\text{norm}(OMS_{map})\geq\alpha\right]$
@@ 3.1.1 Characterisation of the sOMS model
The characterisation of the sOMS has been inspired by the recordings of real cells conducted by Ölveczky et al. [48], where the activity of cells was measured under various conditions: Eye+object, Eye only and Object only. The stimuli consist of moving gratings, with both the background and central region in motion. In the case of Eye+object, both the background and the central region move at different speeds; in Eye only, only the background is moving; and in Object only, only the central part is moving (see Figure 3).
@@
Figure 3: From left to right: Experiment, Parameters, RGB stimuli, Event map and OMS map in three different situations: Eye+object, Eye only and Object only. Where fs{fs} is the spatial frequency in cycles per visual degree (deg) and s{s} is the speed in cycles per frame of the moving grating. The figure shows the sOMS model enhancing object motion, where white and black represent positive and negative polarities on the Event map, and white indicates the spikes of the sOMS model on the sOMS map.
[…]
@@
Figure 4: Average mean firing rate (MFR) and mean inter-spike interval (ISI), along with their respective standard deviations, for the Eye+Object case under two scenarios: (a) varying center ($\sigma_{c}$) and surround ($\sigma_{s}$) values of the kernels (pixels) $\sigma_{c}=1,\sigma_{s}=4$; $\sigma_{c}=2,\sigma_{s}=4$; $\sigma_{c}=3,\sigma_{s}=4$; $\sigma_{c}=4,\sigma_{s}=4$; $\sigma_{c}=2,\sigma_{s}=8$; $\sigma_{c}=4,\sigma_{s}=8$] (Figure 2), and (b) varying both $\sigma$ values and kernel sizes $\sigma_{c}=1,\sigma_{s}=4$, s=8; $\sigma_{c}=4,\sigma_{s}=8$, s=16; $\sigma_{c}=8,\sigma_{s}=16$, s=32].
@@
The Gaussian kernel parameters and sizes were analysed based on their responses in the Eye+Object scenario. Figure 4a) illustrates how decreasing the ratio between $\sigma_{c}$ and $\sigma_{s}$ proportionally reduces both the mean firing rate (MFR) and the mean inter-spike interval (ISI) of the system, due to the diminished inhibitory effect of the surround. These metrics were selected as they provide interpretable and reliable descriptors of spiking activity under varying stimulus conditions. The final two edge cases, in which $\sigma_{s}$ equals the kernel size, show a slight reduction in both metrics, leading to an overall decrease in spiking activity as a result of stronger surround suppression. Larger kernel sizes, as shown in Figure 4 b), confirm the same response trend, with increased activity observed for the largest size. This effect is attributed to the smoothing influence of the Gaussian kernels on the image, resulting in a larger portion of the visual field surpassing the threshold ($\alpha$). An increased MFR would lead to heightened system activity, resulting in unnecessary power consumption, that in neurons chips scales with the number of spikes, as the response is already functional for $\sigma_{c}=1$, $\sigma_{s}=4$, and a kernel size, $s_{k}$, of 8 pixels. This analysis, therefore, validates the parameters proposed in [52].
@@
To demonstrate the scale invariance of the model, the characterisation also includes different spatial frequencies of the stimuli in the Eye+Object scenario. Figure 5 shows an example of the RGB stimuli, the converted event frame, and the OMS maps with different values of spatial frequency, fs{fs}, and the speeds of a cycle on each iteration of the loop, s{s}, for the object (Foreground) and the Background. As shown by the OMS maps in the qualitative assessment in Figure 5, the response of the sOMS model effectively enhances object motion in all cases.
@@
Figure 5: From left to right: Experiment, Parameters, RGB stimuli, Event map and OMS map for the Eye+object case with different spatial frequencies of the background and the foreground. Where fs{fs} is the spatial frequency and s{s} is the speed in cycles per frame of the moving grating. The colors shown on the maps are the same as those in Figure 3.
@@
Further investigations analyze the average MFR and mean ISI of the population activity (see Figure 6) across all experiments presented in Figure 3 and Figure 5, as these metrics provide complementary insights into the dynamics of neural activity. The average MFR captures the overall firing rate, offering a measure of activity intensity, while the mean ISI reflects the temporal structure of spiking, highlighting patterns in neural coordination and timing. The responses confirm the periodicity of the moving gratings, with similar MFR and mean ISI observed across all scenarios, except for Experiment 2 (Eye only). In this case, the OMS model remains effective in suppressing most background activity, even in the absence of object motion; however, the periodicity of the response is lost. This results in a higher mean MFR compared to the other conditions and an elevated mean ISI, indicating a stronger spatial response from the neurons throughout the simulation period, which was not evident in the single example shown in Figure 3. The standard deviation of the mean MFR reflects that only a subset of neurons, corresponding to a specific portion of the visual field, are actively responsive across all cases. In contrast, the standard deviation of the mean ISI becomes extremely large (approaching infinity), as most neurons emit only a single spike during the simulation, resulting in undefined or extremely high interspike intervals.
@@
Figure 6: Average MFR and ISI, along with their respective standard deviations, are reported for all experiments shown in Figure 3 and Figure 5. Experiments 1, 2, and 3 correspond to Eye+Object, Eye only, and Object only, respectively, while experiments 4 through 9 represent Eye+Object scenarios with varying spatial frequencies (fs{fs}) and speeds (s{s}) in cycles per iteration, as depicted in Figure 5.
@@
Figure 7: Example of OMS map for different speeds (where $s_{F}\leq s_{B}$) for the background ($s_{B}$) and foreground ($s_{F}$) in the case of the Eye+object scenario for the Background (fsfs=0.3) and the Foreground (fsfs=3).
In all cases examined within Experiments 10–16 (Figure 7), the sOMS model appears unable to enhance object motion effectively (same applies for $\tau$ = 0.01), resulting in poor performance in object motion segmentation when egomotion-induced events are generated from a background moving faster than the foreground. The failure of motion detection can be attributed to two critical mechanisms. First, spatiotemporal integration between center and surround regions becomes ineffective when object-induced events are temporally sparse or significantly lower than background-induced events. Under such conditions, the membrane potential fails to reach the threshold required for motion detection or the dominance of background activity suppresses the differential signal needed to distinguish motion.
Second, from an active vision perspective, the motion dynamics of the visual sensor critically impact detection reliability. The challenge of motion detection under rapid camera movement is a well-documented issue in optic flow estimation [83]. When camera movement exceeds a specific velocity threshold, local motion contrast becomes unresolvable, directly compromising the fundamental assumptions of optical flow estimation.
@@ 3.2 SNN Proto-Object model
This implementation builds upon the event-driven, saliency-based proto-object model previously developed for GPU [30] and SpiNNaker [39]. The primary difference from the earlier implementations is the use of the LIF sCNN sinabs layer, with weights represented by the von Mises kernel distribution (Equation 2).
$VM_{\theta}(x,y)=\frac{\exp(\rho\cdot R_{0}\cdot\cos(\tan^{-1}(-y,x)-\theta)}{I_{0}(\sqrt{x^{2}+y^{2}-R_{0}})}$ (2)
@@
$G_{1}=\sum_{\theta}\sum_{s}B1_{\theta,s}\cdot VM_{\theta,s}$ $G_{2}=\sum_{\theta}\sum_{s}B2_{\theta,s}\cdot VM_{\theta+\pi,s}$ $G_{1}^{*}=\sum_{\theta}\sum_{s}B1_{\theta,s}\cdot VM_{\theta+\pi,s}$ $G_{2}^{*}=\sum_{\theta}\sum_{s}B2_{\theta,s}\cdot VM_{\theta,s}$ (4)
These orientations and scales ($s=3$) are pooled (Equation 4) to produce the final grouping response, which generates the saliency map of the scene and identifies the $(x,y)$ coordinates of the most salient point (Equation 5 and Figure 1).
$P(x,y)=\arg\max_{(x,y)}\left[(G_{1}-G_{1}^{*})+(G_{2}-G_{2}^{*})\right]$ (5)